OpenStudy (anonymous):

On June 1, Mary Barkley deposited $1,000 in a savings account that pays 5.5% interest compounded daily. On June 26, how much interest has she earned on principal?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

hi

OpenStudy (loser66):

hi

OpenStudy (anonymous):

any idea how to solve this any pointers?

OpenStudy (loser66):

you have formula, right? just apply

OpenStudy (anonymous):

I have. I dont think I am applying my numbers right

OpenStudy (loser66):

show me your work

OpenStudy (anonymous):

A=P(1+r)^n here is the formula

OpenStudy (anonymous):

the pay periods and rate are giving me trouble . dont know if im supposed to be using 25 days (june 1 - june 26) or 365 days. either way, the numbers never come out right

OpenStudy (loser66):

I think 25 days

OpenStudy (anonymous):

ya but like i said, even with those numbers it doesnt come out

OpenStudy (loser66):

type all

OpenStudy (loser66):

I mean your work

OpenStudy (anonymous):

Rate = .055/25 Rate = .001964285 A=P(1.00 + .001964285)^n A=P(1.001964285)^n A=P(1.001964285)^25 A=P(1.050282271 A=1000(1.050282271 A=1050.282271

OpenStudy (anonymous):

i tried multiplying like you said and i got an error when i got to the exponent caluclation

OpenStudy (loser66):

I see you apply monthly, right?

OpenStudy (anonymous):

what do you mean

OpenStudy (loser66):

http://www.purplemath.com/modules/expofcns4.htm because I don't think you apply the correct formula

OpenStudy (anonymous):

that was a secondary formula that my teacher gave me. the one im using now is the one she recommended, but I guess ill give it a shot